Question 5 of 10 (1 point) View problem in a pop-up A missile guidance system has seven fail-safe components. The probability of each failing is 0.1. Assume the variable is binomial. Find the following probabilities. Round your answers to three decimal places 5.3 Section Exercise 16 Part 1 a. Exactly three will fail P (exactly three will fail)-.023 Part 2 b. More than three will fail. P (more than three will fail003 Part 3 out of 4 c. All will fail. P (all will fail) = L
Answer #1

for binomial distribution ; parameter n=7 and p=0.1

c)Probability all will fail =P(X=7) =7C7 (0.1)7*(0.9)0 =1*0.0000001*1 =~0.000